Two bags of M&Ms were emptied into a coffee can. When you sort and count the colors and tyles of M&Ms, you find the following quantities.
Plain Peanut
Red. 11. 9
Orange. 8. 3
Yellow. 9. 7
Green. 5. 10
Blue. 12. 14
Brown. 6. 4
When you aelect one candie at random, what is the probability that is a blue one or a plain one?
Answer #1

Based on the above data,

When you select one candie at random, the probability that is a blue one or a plain one = (12+14+11+8+9+5+6)/(11+9+8+3+9+7+5+10+12+14+6+4) = 65/98 = 0.6633